What information does the payback period provide?

Suppose ABC Telecom Inc.’s CFO is evaluating a project with the following cash inflows. She does not know the project’s initial cost; however, she does know that the project’s regular payback period is 2.5 years.

Year

Cash Flow

Year 1 $375,000
Year 2 $500,000
Year 3 $425,000
Year 4 $475,000

If the project’s weighted average cost of capital (WACC) is 7%, what is its NPV?

$470,336

$408,988

$490,786

$368,089

Which of the following statements indicate a disadvantage of using the discounted payback period for capital budgeting decisions? Check all that apply.

The discounted payback period does not take the time value of money into account.

The discounted payback period does not take the project’s entire life into account.

The discounted payback period is calculated using net income instead of cash flows.
Answer #1

& Initial Payback = 2-5 years Investment = 375000 + 500000 + 425000 - 1087500 2 MPV = 375000 + 500000 1.07 1,072 & 425000 + 4